4 of 5 Problem 10: (30 Points) t = -0...+00. In the following circuit calculate v. (t) Answer: v.() = 8 V + 20% vo(t)
Answer #1

→ 2H Solution for watco, Switch is closed and the equivalent circuit is drawn as, il-a)=0, i sor) is found as steady stati sv4,1) +(1+) { t) Substitute V(!)-12 + 1/4 = Ut) Volt) - 6 - (1) I Volt) = 6+ 2042 for thol. for t>o, Switch is open, the equ